# Efficient Epistemic Uncertainty Estimation in Regression Ensemble Models Using Pairwise-Distance Estimators

### v3

@@ Appendix D Additional Results
Env Acq. Batch Random BatchBALD BADGE BAIT LCMD Monte Carlo KL Bhatt
hetero 10 0.54 $\pm$ 0.27 -0.59 $\pm$ 0.55 -0.49 $\pm$ 0.61 -0.65 $\pm$ 0.54 -0.43 $\pm$ 0.26 0.95 $\pm$ 0.12 0.93 $\pm$ 0.14 0.99 $\pm$ 0.12
25 0.79 $\pm$ 0.18 -0.57 $\pm$ 0.35 -0.53 $\pm$ 0.24 -0.64 $\pm$ 0.31 -0.32 $\pm$ 0.21 1.06 $\pm$ 0.06 1.06 $\pm$ 0.12 1.11 $\pm$ 0.09
50 0.8 $\pm$ 0.18 -0.25 $\pm$ 0.2 -0.29 $\pm$ 0.22 -0.36 $\pm$ 0.16 -0.1 $\pm$ 0.27 1.08 $\pm$ 0.04 1.07 $\pm$ 0.16 1.15 $\pm$ 0.14
100 0.86 $\pm$ 0.14 -0.05 $\pm$ 0.44 -0.06 $\pm$ 0.42 -0.31 $\pm$ 0.15 0.06 $\pm$ 0.21 1.09 $\pm$ 0.1 1.12 $\pm$ 0.1 1.14 $\pm$ 0.13
bimodal 10 -30.57 $\pm$ 12.61 0.68 $\pm$ 0.29 0.7 $\pm$ 0.48 0.75 $\pm$ 0.22 0.74 $\pm$ 0.39 1.08 $\pm$ 0.09 1.11 $\pm$ 0.1 1.11 $\pm$ 0.1
25 -16.55 $\pm$ 9.04 0.74 $\pm$ 0.26 0.8 $\pm$ 0.23 0.67 $\pm$ 0.32 0.81 $\pm$ 0.13 1.21 $\pm$ 0.09 1.17 $\pm$ 0.1 1.27 $\pm$ 0.09
50 -10.85 $\pm$ 8.5 0.67 $\pm$ 0.28 0.84 $\pm$ 0.22 0.82 $\pm$ 0.24 0.74 $\pm$ 0.17 1.22 $\pm$ 0.11 1.2 $\pm$ 0.1 1.21 $\pm$ 0.11
100 -6.19 $\pm$ 8.66 0.84 $\pm$ 0.25 0.83 $\pm$ 0.25 0.92 $\pm$ 0.33 0.89 $\pm$ 0.27 1.26 $\pm$ 0.13 1.26 $\pm$ 0.1 1.26 $\pm$ 0.14
Pendulum 10 -5.55 $\pm$ 6.8 -0.99 $\pm$ 2.71 -8.64 $\pm$ 4.52 -0.2 $\pm$ 3.58 0.12 $\pm$ 3.42 5.81 $\pm$ 3.29 6.17 $\pm$ 1.86 5.8 $\pm$ 1.41
25 -0.78 $\pm$ 5.92 5.25 $\pm$ 2.42 5.52 $\pm$ 1.81 6.94 $\pm$ 1.58 7.09 $\pm$ 1.69 8.72 $\pm$ 0.45 8.45 $\pm$ 0.55 8.3 $\pm$ 0.7
50 3.77 $\pm$ 1.8 8.41 $\pm$ 1.55 7.8 $\pm$ 1.71 8.24 $\pm$ 2.06 8.38 $\pm$ 1.51 9.77 $\pm$ 0.36 9.95 $\pm$ 0.5 9.59 $\pm$ 0.27
100 5.19 $\pm$ 2.45 9.46 $\pm$ 1.03 8.94 $\pm$ 1.15 8.9 $\pm$ 1.36 10.22 $\pm$ 1.14 10.16 $\pm$ 0.7 10.49 $\pm$ 0.23 9.58 $\pm$ 0.8
Hopper 10 -136.85 $\pm$ 53.59 1.49 $\pm$ 6.05 -3.93 $\pm$ 6.13 0.33 $\pm$ 6.2 4.69 $\pm$ 2.87 13.94 $\pm$ 5.54 11.59 $\pm$ 4.32 11.44 $\pm$ 2.31
25 -28.0 $\pm$ 28.16 15.86 $\pm$ 2.43 13.64 $\pm$ 2.93 17.54 $\pm$ 2.08 18.53 $\pm$ 2.05 23.18 $\pm$ 1.41 22.71 $\pm$ 0.86 22.29 $\pm$ 0.95
50 -0.55 $\pm$ 6.82 22.46 $\pm$ 2.42 21.98 $\pm$ 3.26 23.06 $\pm$ 3.95 23.66 $\pm$ 3.18 26.43 $\pm$ 1.37 25.72 $\pm$ 1.04 25.0 $\pm$ 1.6
100 8.67 $\pm$ 3.39 25.9 $\pm$ 1.89 24.4 $\pm$ 2.98 23.82 $\pm$ 2.39 26.07 $\pm$ 1.64 26.62 $\pm$ 1.14 26.96 $\pm$ 1.11 26.88 $\pm$ 0.91
Ant 10 -0.5 $\pm$ 16.68 1.09 $\pm$ 12.41 -2.24 $\pm$ 12.36 -1.35 $\pm$ 13.33 17.54 $\pm$ 6.36 24.48 $\pm$ 5.88 25.41 $\pm$ 5.08 21.59 $\pm$ 9.0
25 31.08 $\pm$ 5.4 22.52 $\pm$ 4.73 16.87 $\pm$ 7.39 22.0 $\pm$ 4.7 23.35 $\pm$ 4.48 39.77 $\pm$ 3.34 40.62 $\pm$ 2.85 37.58 $\pm$ 2.37
50 43.37 $\pm$ 1.85 28.03 $\pm$ 3.53 26.73 $\pm$ 3.35 30.84 $\pm$ 2.19 28.92 $\pm$ 2.59 45.67 $\pm$ 1.76 45.07 $\pm$ 1.62 44.77 $\pm$ 1.39
100 47.52 $\pm$ 1.36 33.69 $\pm$ 3.1 31.57 $\pm$ 3.42 35.39 $\pm$ 2.6 31.89 $\pm$ 3.94 49.29 $\pm$ 0.78 48.12 $\pm$ 2.05 46.52 $\pm$ 2.42
Humanoid 10 -200.61 $\pm$ 20.02 -153.76 $\pm$ 48.74 -200.56 $\pm$ 37.2 -230.61 $\pm$ 9.92 -220.73 $\pm$ 20.19 -249.1 $\pm$ 0.69 -221.09 $\pm$ 10.68 -201.28 $\pm$ 22.78
25 -223.54 $\pm$ 13.72 -231.51 $\pm$ 16.72 -238.87 $\pm$ 14.85 -245.86 $\pm$ 1.8 -226.07 $\pm$ 26.57 -247.84 $\pm$ 1.74 -246.29 $\pm$ 1.92 -247.16 $\pm$ 1.04
50 -244.15 $\pm$ 1.52 -247.03 $\pm$ 2.56 -246.62 $\pm$ 4.16 -248.41 $\pm$ 0.93 -235.55 $\pm$ 13.35 -247.76 $\pm$ 2.11 -249.39 $\pm$ 0.43 -249.57 $\pm$ 0.62
100 -247.82 $\pm$ 1.28 -249.38 $\pm$ 0.69 -249.15 $\pm$ 0.69 -249.71 $\pm$ 0.32 -241.1 $\pm$ 16.55 -246.47 $\pm$ 2.99 -249.76 $\pm$ 0.4 -249.97 $\pm$ 0.05
@@ Appendix E Probabilistic Network Ensembles
In addition to ensembles of normalizing flows, we also explored the use of probabilistic network ensembles (PNEs) as a common approach for capturing uncertainty (Chua et al., 2018; Kurutach et al., 2018; Lakshminarayanan et al., 2017). The PNEs were constructed by employing fixed dropout masks, where each ensemble component modeled a Gaussian distribution. The models were trained using negative log likelihood, with weights randomly initialized and bootstrapped samples from the training set to create diversity amongst the componenents. Our findings paralleled those of Nflows Base, in that PaiDEs performed similarly or better than baselines and statistically significantly in higher dimensions. These results are presented in s Figure 11 and Figure 12, as well as Table 4 and Table 5. Note that PNEs did not perform as well as Nflows Base as they are not as expressive this can be seen for the bimodal setting in Figure 10. Furthermore, we have included the 1D graphs illustrating the performance of PNEs in hetero and bimodal in Figure 10. Similarly to before, we also provide a comparison of the MC estimator as the number of samples increased in Figure 13. While we could have implemented ensembles with different output distributions to better fit the data, we decided to stick with Gaussians as they seem to be the most frequent choice (Chua et al., 2018; Kurutach et al., 2018; Lakshminarayanan et al., 2017). Moreover, in order to pick the best distributional fit would require a hyper-parameter search or apriori knowledge whereas NFs will learn the best distributional fit.
@@
Env Acq. Batch Random BatchBALD BADGE BAIT LCMD Monte Carlo KL Bhatt
hetero 10 3.9 $\pm$ 1.47 6.9 $\pm$ 1.15 6.91 $\pm$ 1.77 6.9 $\pm$ 1.58 6.49 $\pm$ 1.63 1.94 $\pm$ 0.34 2.11 $\pm$ 0.34 2.36 $\pm$ 0.75
25 4.42 $\pm$ 1.43 4.43 $\pm$ 2.81 4.06 $\pm$ 2.58 5.02 $\pm$ 2.91 3.85 $\pm$ 2.51 1.75 $\pm$ 0.27 1.86 $\pm$ 0.39 1.96 $\pm$ 0.44
50 3.96 $\pm$ 1.71 4.06 $\pm$ 2.64 5.12 $\pm$ 2.78 5.22 $\pm$ 2.93 4.87 $\pm$ 2.36 1.78 $\pm$ 0.51 1.68 $\pm$ 0.34 1.67 $\pm$ 0.27
100 4.28 $\pm$ 1.45 5.1 $\pm$ 2.92 4.54 $\pm$ 2.89 6.01 $\pm$ 2.82 3.85 $\pm$ 2.5 1.66 $\pm$ 0.31 1.83 $\pm$ 0.51 1.98 $\pm$ 0.81
bimodal 10 6.52 $\pm$ 0.88 6.05 $\pm$ 0.07 6.19 $\pm$ 0.27 6.11 $\pm$ 0.15 6.05 $\pm$ 0.06 6.01 $\pm$ 0.04 6.02 $\pm$ 0.04 6.01 $\pm$ 0.04
25 6.1 $\pm$ 0.08 6.26 $\pm$ 0.75 6.31 $\pm$ 0.83 6.3 $\pm$ 0.88 6.28 $\pm$ 0.8 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04
50 6.13 $\pm$ 0.13 6.01 $\pm$ 0.04 6.04 $\pm$ 0.07 6.02 $\pm$ 0.03 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04
100 6.06 $\pm$ 0.08 6.01 $\pm$ 0.03 6.03 $\pm$ 0.03 6.01 $\pm$ 0.04 6.01 $\pm$ 0.03 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04
Pendulum 10 0.31 $\pm$ 0.11 0.15 $\pm$ 0.02 0.16 $\pm$ 0.05 0.15 $\pm$ 0.03 0.16 $\pm$ 0.03 0.19 $\pm$ 0.04 0.18 $\pm$ 0.06 0.2 $\pm$ 0.07
25 0.27 $\pm$ 0.09 0.11 $\pm$ 0.03 0.12 $\pm$ 0.03 0.11 $\pm$ 0.03 0.12 $\pm$ 0.03 0.16 $\pm$ 0.02 0.12 $\pm$ 0.04 0.13 $\pm$ 0.04
50 0.25 $\pm$ 0.07 0.08 $\pm$ 0.02 0.09 $\pm$ 0.02 0.08 $\pm$ 0.02 0.09 $\pm$ 0.03 0.15 $\pm$ 0.03 0.08 $\pm$ 0.02 0.09 $\pm$ 0.03
100 0.28 $\pm$ 0.06 0.07 $\pm$ 0.03 0.07 $\pm$ 0.02 0.08 $\pm$ 0.02 0.07 $\pm$ 0.02 0.16 $\pm$ 0.03 0.09 $\pm$ 0.06 0.08 $\pm$ 0.05
Hopper 10 1.85 $\pm$ 0.14 1.44 $\pm$ 0.31 1.12 $\pm$ 0.29 1.16 $\pm$ 0.37 1.09 $\pm$ 0.36 0.94 $\pm$ 0.16 1.19 $\pm$ 0.21 1.14 $\pm$ 0.18
25 1.9 $\pm$ 0.14 0.6 $\pm$ 0.13 0.66 $\pm$ 0.19 0.55 $\pm$ 0.12 0.54 $\pm$ 0.14 0.73 $\pm$ 0.13 0.72 $\pm$ 0.08 0.71 $\pm$ 0.1
50 1.83 $\pm$ 0.13 0.54 $\pm$ 0.15 0.55 $\pm$ 0.15 0.51 $\pm$ 0.12 0.51 $\pm$ 0.12 0.57 $\pm$ 0.06 0.57 $\pm$ 0.06 0.57 $\pm$ 0.05
100 1.72 $\pm$ 0.14 0.4 $\pm$ 0.05 0.49 $\pm$ 0.13 0.43 $\pm$ 0.05 0.47 $\pm$ 0.08 0.5 $\pm$ 0.04 0.44 $\pm$ 0.04 0.49 $\pm$ 0.05
Ant 10 1.89 $\pm$ 0.12 2.27 $\pm$ 0.13 2.24 $\pm$ 0.12 2.15 $\pm$ 0.18 2.09 $\pm$ 0.19 1.87 $\pm$ 0.25 1.62 $\pm$ 0.16 1.52 $\pm$ 0.18
25 1.74 $\pm$ 0.13 2.15 $\pm$ 0.21 2.14 $\pm$ 0.18 2.12 $\pm$ 0.25 2.08 $\pm$ 0.13 1.67 $\pm$ 0.25 1.43 $\pm$ 0.12 1.39 $\pm$ 0.12
50 1.55 $\pm$ 0.1 2.2 $\pm$ 0.13 2.16 $\pm$ 0.14 2.11 $\pm$ 0.18 2.07 $\pm$ 0.16 1.61 $\pm$ 0.22 1.29 $\pm$ 0.06 1.3 $\pm$ 0.04
100 1.5 $\pm$ 0.12 2.01 $\pm$ 0.2 1.99 $\pm$ 0.2 1.97 $\pm$ 0.15 2.02 $\pm$ 0.16 1.73 $\pm$ 0.22 1.26 $\pm$ 0.08 1.24 $\pm$ 0.04
Humanoid 10 10.73 $\pm$ 0.32 11.63 $\pm$ 0.19 11.22 $\pm$ 0.29 10.95 $\pm$ 0.17 10.98 $\pm$ 0.16 10.41 $\pm$ 0.53 9.95 $\pm$ 1.1 10.65 $\pm$ 0.84
25 10.38 $\pm$ 0.22 11.36 $\pm$ 0.26 11.07 $\pm$ 0.26 10.91 $\pm$ 0.25 11.06 $\pm$ 0.23 10.44 $\pm$ 0.55 6.54 $\pm$ 2.13 8.22 $\pm$ 2.47
50 10.29 $\pm$ 0.19 11.37 $\pm$ 0.3 10.98 $\pm$ 0.22 10.9 $\pm$ 0.25 11.01 $\pm$ 0.22 10.34 $\pm$ 0.81 4.91 $\pm$ 0.24 6.33 $\pm$ 2.24
100 10.17 $\pm$ 0.3 11.32 $\pm$ 0.16 10.99 $\pm$ 0.26 11.05 $\pm$ 0.21 11.2 $\pm$ 0.23 10.57 $\pm$ 0.41 4.84 $\pm$ 0.24 6.25 $\pm$ 2.32
@@
Figure 14: The same plots as Figure 2 except there is no mapping applied and the corresponding estimates for $I(y,W)$ are shown the right y-axes for the right plots.

### v4

@@ Appendix D Additional Results
Env Acq. Batch Random BatchBALD BADGE BAIT MC (BALD) KL (ours) Bhatt (ours)
hetero 10 0.54 $\pm$ 0.27 0.21 $\pm$ 1.04 -0.67 $\pm$ 2.48 0.27 $\pm$ 0.7 0.95 $\pm$ 0.12 0.93 $\pm$ 0.14 $9±\pm 0$.12
[…]
@@ Appendix E Probabilistic Network Ensembles
In addition to ensembles of normalizing flows, we also explored the use of probabilistic network ensembles (PNEs) as a common approach for capturing uncertainty [12, 38, 39]. The PNEs were constructed by employing fixed dropout masks, where each ensemble component modeled a Gaussian distribution. The models were trained using negative log likelihood, with weights randomly initialized and bootstrapped samples from the training set to create diversity amongst the componenents. Our findings paralleled those of Nflows Base, in that PairEpEsts performed similarly or better than baselines and statistically significantly in higher dimensions. These results are presented in Figure 13 and Figure 14, as well as Table 9 and Table 10. Note that PNEs did not perform as well as Nflows Base as they are not as expressive this can be seen for the bimodal setting in Figure 12. Furthermore, we have included the 1D graphs illustrating the performance of PNEs in hetero and bimodal in Figure 12. Similarly to before, we also provide a comparison of the MC estimator as the number of samples increased in Figure 15. While we could have implemented ensembles with different output distributions to better fit the data, we decided to stick with Gaussians as they seem to be the most frequent choice [12, 38, 39]. Moreover, in order to pick the best distributional fit would require a hyper-parameter search or apriori knowledge whereas NFs will learn the best distributional fit.
@@
Env Acq. Batch Random BatchBALD BADGE BAIT MC (BALD) KL (ours) Bhatt (ours)
hetero 10 3.9 $\pm$ 1.47 6.9 $\pm$ 1.15 6.91 $\pm$ 1.77 6.9 $\pm$ 1.58 1.94 $\pm$ 0.34 2.11 $\pm$ 0.34 2.36 $\pm$ 0.75
25 4.42 $\pm$ 1.43 4.43 $\pm$ 2.81 4.06 $\pm$ 2.58 5.02 $\pm$ 2.91 1.75 $\pm$ 0.27 1.86 $\pm$ 0.39 1.96 $\pm$ 0.44
50 3.96 $\pm$ 1.71 4.06 $\pm$ 2.64 5.12 $\pm$ 2.78 5.22 $\pm$ 2.93 1.78 $\pm$ 0.51 1.68 $\pm$ 0.34 1.67 $\pm$ 0.27
100 4.28 $\pm$ 1.45 5.1 $\pm$ 2.92 4.54 $\pm$ 2.89 6.01 $\pm$ 2.82 1.66 $\pm$ 0.31 1.83 $\pm$ 0.51 1.98 $\pm$ 0.81
bimodal 10 6.52 $\pm$ 0.88 6.05 $\pm$ 0.07 6.19 $\pm$ 0.27 6.11 $\pm$ 0.15 6.01 $\pm$ 0.04 6.02 $\pm$ 0.04 6.01 $\pm$ 0.04
25 6.1 $\pm$ 0.08 6.26 $\pm$ 0.75 6.31 $\pm$ 0.83 6.3 $\pm$ 0.88 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04
50 6.13 $\pm$ 0.13 6.04 $\pm$ 0.07 6.02 $\pm$ 0.03 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04
100 6.06 $\pm$ 0.08 6.01 $\pm$ 0.03 6.03 $\pm$ 0.03 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04
Pendulum 10 0.31 $\pm$ 0.11 0.15 $\pm$ 0.02 0.16 $\pm$ 0.05 0.15 $\pm$ 0.03 0.19 $\pm$ 0.04 0.18 $\pm$ 0.06 0.2 $\pm$ 0.07
25 0.27 $\pm$ 0.09 0.11 $\pm$ 0.03 0.12 $\pm$ 0.03 0.11 $\pm$ 0.03 0.16 $\pm$ 0.02 0.12 $\pm$ 0.04 0.13 $\pm$ 0.04
50 0.25 $\pm$ 0.07 0.08 $\pm$ 0.02 0.09 $\pm$ 0.02 0.08 $\pm$ 0.02 0.15 $\pm$ 0.03 0.08 $\pm$ 0.02 0.09 $\pm$ 0.03
100 0.28 $\pm$ 0.06 0.07 $\pm$ 0.03 0.07 $\pm$ 0.02 0.08 $\pm$ 0.02 0.16 $\pm$ 0.03 0.09 $\pm$ 0.06 0.08 $\pm$ 0.05
Hopper 10 1.85 $\pm$ 0.14 1.44 $\pm$ 0.31 1.12 $\pm$ 0.29 1.16 $\pm$ 0.37 0.94 $\pm$ 0.16 1.19 $\pm$ 0.21 1.14 $\pm$ 0.18
25 1.9 $\pm$ 0.14 0.6 $\pm$ 0.13 0.66 $\pm$ 0.19 $5±\pm 0$.12 0.73 $\pm$ 0.13 0.72 $\pm$ 0.08 0.71 $\pm$ 0.1
50 1.83 $\pm$ 0.13 0.54 $\pm$ 0.15 0.55 $\pm$ 0.15 0.51 $\pm$ 0.12 0.57 $\pm$ 0.06 0.57 $\pm$ 0.06 0.57 $\pm$ 0.05
100 1.72 $\pm$ 0.14 0.4 $\pm$ 0.05 0.49 $\pm$ 0.13 0.43 $\pm$ 0.05 0.5 $\pm$ 0.04 0.44 $\pm$ 0.04 0.49 $\pm$ 0.05
Ant 10 1.89 $\pm$ 0.12 2.27 $\pm$ 0.13 2.24 $\pm$ 0.12 2.15 $\pm$ 0.18 1.87 $\pm$ 0.25 1.62 $\pm$ 0.16 1.52 $\pm$ 0.18
25 1.74 $\pm$ 0.13 2.15 $\pm$ 0.21 2.14 $\pm$ 0.18 2.12 $\pm$ 0.25 1.67 $\pm$ 0.25 1.43 $\pm$ 0.12 1.39 $\pm$ 0.12
50 1.55 $\pm$ 0.1 2.2 $\pm$ 0.13 2.16 $\pm$ 0.14 2.11 $\pm$ 0.18 1.61 $\pm$ 0.22 1.29 $\pm$ 0.06 1.3 $\pm$ 0.04
100 1.5 $\pm$ 0.12 2.01 $\pm$ 0.2 1.99 $\pm$ 0.2 1.97 $\pm$ 0.15 1.73 $\pm$ 0.22 1.26 $\pm$ 0.08 1.24 $\pm$ 0.04
Humanoid 10 10.73 $\pm$ 0.32 11.63 $\pm$ 0.19 11.22 $\pm$ 0.29 10.95 $\pm$ 0.17 10.41 $\pm$ 0.53 9.95 $\pm$ 1.1 10.65 $\pm$ 0.84
25 10.38 $\pm$ 0.22 11.36 $\pm$ 0.26 11.07 $\pm$ 0.26 10.91 $\pm$ 0.25 10.44 $\pm$ 0.55 6.54 $\pm$ 2.13 8.22 $\pm$ 2.47
50 10.29 $\pm$ 0.19 11.37 $\pm$ 0.3 10.98 $\pm$ 0.22 10.9 $\pm$ 0.25 10.34 $\pm$ 0.81 4.91 $\pm$ 0.24 6.33 $\pm$ 2.24
100 10.17 $\pm$ 0.3 11.32 $\pm$ 0.16 10.99 $\pm$ 0.26 11.05 $\pm$ 0.21 10.57 $\pm$ 0.41 4.84 $\pm$ 0.24 6.25 $\pm$ 2.32
@@
Figure 16: The bias introduced by PairEpEsts for PNEs compared to the MC method on the 1D environments.